A model rocket has a horizontal range of 260 m on
Answer #1

horizontal range on level ground R = v^2 sin(2@) / g

260 = v^2 sin90 / 9.81

v = 50.50 m/s .....max velocity of rocket.

R = (v^2 / g cos^225 ) [ sin(90+25) + sin25 ]

R = ( 50.50^2 / (9.81 x 0.906^2)) [ sin115 + sin25]

R = 393.28 m this is along the incline.

in horizontal = Rcos25 = 356.43m